# Localized RETE for Incremental Graph Queries with Nested Graph Conditions

### v1

@@ Appendix A Technical Details
For each nested query with the form $(Q^{\prime},\exists(a:Q^{\prime}\rightarrow Q^{\prime\prime},\psi^{\prime}))$, $N^{sat}$ contains the subnets $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$ and $(N^{\Phi}_{Q^{\prime\prime}},p^{\Phi}_{Q^{\prime\prime}})=localize(Q^{\prime\prime})$, a marking-sensitive union node $[\cup]^{\Phi}$, a marking-sensitive semi-join $[\ltimes]^{\Phi}$, as well as the request projection structure $RPS_{l}=RPS([\cup]^{\Phi},N^{\Phi}_{Q^{\prime}})$. $N$ contains a RETE subnet $(N_{Q^{\prime}},p_{Q^{\prime}})$ computing matches for $Q^{\prime}$ and a semi-join []delimited-[]left-normal-factor-semidirect-product[$\ltimes$][ ⋉ ].
By Theorem 5 and because the additional dependency on $RPS_{l}$ does not affect the worst case of a fully populated RETE net $(N^{\Phi}_{Q},p^{\Phi}_{Q})$, it holds that $\sum_{n^{\Phi}_{Q^{\prime}}\in V^{N^{\Phi}_{Q^{\prime}}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q^{\prime}})}|m|\leq 7\cdot|\mathcal{C}|_{N_{Q^{\prime}}}|_{e}$. Also, by the semantics of the marking-sensitive semi-join and because the result set for $p^{\Phi}_{Q^{\prime}}$ contains each match for $Q^{\prime}$ at most once by Lemma 12 in [BG24b], it must hold that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\ltimes]^{\Phi})}|m|\leq\sum_{m\in\mathcal{C}(p_{Q^{\prime}})}|m|\leq|\mathcal{C}|_{N_{Q^{\prime}}}|_{e}$.
Any nested query with the form $(Q^{\prime},\exists(a:Q^{\prime}\rightarrow Q^{\prime\prime},\psi^{\prime}))$ must have some nested query $(Q^{\prime\prime},\exists(a:Q^{\prime\prime}\rightarrow Q^{\prime\prime\prime},\psi^{\prime\prime}))$ or $(Q^{\prime\prime},\texttt{true})$ in order for $\psi$ by the definition of nested graph conditions. In either case, $N$ has to contain a corresponding subnet $(N_{Q^{\prime\prime}},p_{Q^{\prime\prime}})$ computing matches for $Q^{\prime\prime}$. It then holds for this subnet $N_{Q^{\prime\prime}}$ by Theorem 5 that $\sum_{n^{\Phi}_{Q^{\prime\prime}}\in V^{N^{\Phi}_{Q^{\prime\prime}}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q^{\prime\prime}})}|m|\leq 7\cdot|\mathcal{C}|_{N_{Q^{\prime\prime}}}|_{e}$. It also must hold that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\cup]^{\Phi})}|m|\leq\sum_{m\in\mathcal{C}(p_{Q^{\prime\prime}})}|m|\leq|\mathcal{C}|_{N_{Q^{\prime\prime}}}|_{e}$ and $\sum_{n^{\Phi}\in V^{RPS_{l}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 2\cdot|\mathcal{C}|_{N_{Q^{\prime\prime}}}|_{e}$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ for host graph $H$ and relevant subgraph $H_{p}$ corresponding to $(N,p)$. Assuming that storing a match $m$ requires an amount of memory in $O(|m|)$ and storing an element from $\overline{\mathbb{N}}$ requires an amount of memory in $O(1)$, storing $\mathcal{C}^{\Phi}$ requires memory in $O(|\mathcal{C}|_{e})$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and and $\mathcal{C}^{\Phi}_{0}$ the empty configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ corresponding to $(N,p)$. Executing $(N^{\Phi},p^{\Phi})$ via $execute(order^{\Psi}(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(T\cdot|Q_{max}|)$ steps, with $T=\sum_{n\in V^{N}}|\mathcal{C}(n)|$ and $|Q_{max}|$ the size of the largest graph among $Q$ and graphs in $\psi$.
@@ A.5. Supplementary Lemmata
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net with $Q$ the associated query graph of $p$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\cdot\sum_{n\in V^{N}\setminus\{p\}}\sum_{m\in\mathcal{C}(n)}|m|+5\cdot\sum_{m\in\mathcal{C}(p)}|m|$.
@@
Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net with $Q$ the associated query graph of $p$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\sum_{n\in V^{N}\setminus\{p\}}|\mathcal{C}(n)|+7\times|\mathcal{C}(p)|$.
@@
Let $H$ be a graph, $(Q,\psi)$ a graph query, $(X^{\Phi},p_{X}^{\Phi})$ a modular extension of the localized RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\phi)$, and $\mathcal{C}$ a configuration for $(X^{\Phi},p_{X}^{\Phi})$ that is consistent for all nodes in $V^{N^{\Phi}}$. Furthermore, let $x\in\chi(N^{\Phi})$, $Q_{v}=(\{v\},\emptyset,\emptyset,\emptyset)$ the query subgraph associated with $x$, and $(m^{\prime},\infty)\in\mathcal{C}(x)$ for some match $m^{\prime}:Q_{v}\rightarrow H$. It then holds that $\forall(m,\phi)\in\mathcal{C}(p^{\Phi}):m\models\psi\wedge\forall m\in\mathcal{M}^{Q}_{H}:m(v)=m^{\prime}(v)\wedge m\models\psi\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi})$.
@@
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $N^{\Phi}$ consists of the localized RETE net for $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, a request projection structure $RPS_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q^{\prime},\psi^{\prime})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q^{\prime},\psi^{\prime})}$.
@@
By Lemma 1 in the preprint version of our conference paper [BG24b], it immediately follows that $\forall m\in\mathcal{M}^{Q}_{H}:m(v)=m^{\prime}(v)\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q})$. Assuming that the Lemma holds for the localized RETE net $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, by the definition of $RPS_{l}$ it must also hold that $\forall(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q}),m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m=m^{\prime}\circ a\wedge m^{\prime}\models\psi^{\prime}\Rightarrow(m^{\prime},\infty)\in\mathcal{C}(p^{\Phi}_{(Q^{\prime},\psi^{\prime})})$ and $\forall(m^{\prime},\phi^{\prime})\in\mathcal{C}(p^{\Phi}_{(Q^{\prime},\psi^{\prime})}):m^{\prime}\models\psi^{\prime}$. From the semantics of the marking-sensitive semi-join then follows the satisfaction of the Lemma.
@@
For a nested condition of the form $\psi=\neg\psi^{\prime}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, a request projection structure $RPS_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi^{\prime})})$, and a marking-sensitive anti-join node $p^{\Phi}=[\rhd]^{\Phi}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q,\psi^{\prime})}$.
@@
By Lemma 1 in the preprint version of our conference paper [BG24b], it immediately follows that $\forall m\in\mathcal{M}^{Q}_{H}:m(v)=m^{\prime}(v)\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q})$. Assuming that the Lemma holds for the localized RETE net $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, by the definition of $RPS_{l}$ it must also hold that $\forall(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q}):m\models\psi^{\prime}\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{(Q,\psi^{\prime})})$ and $\forall(m,\phi)\in\mathcal{C}(p^{\Phi}_{(Q,\psi^{\prime})}):m\models\psi^{\prime}$. From the semantics of the marking-sensitive anti-join then follows the satisfaction of the Lemma.
@@
For a nested condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, a request projection structure $RPS_{1}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi_{1})})$, a marking-sensitive semi-join node $[\ltimes]^{\Phi}_{1}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q^{\prime},\psi^{\prime})}$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, a request projection structure $RPS_{2}=RPS([\ltimes]^{\Phi}_{1},N^{\Phi}_{(Q,\psi_{2})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}_{2}$ with left dependency $[\ltimes]^{\Phi}_{1}$ and right dependency $p^{\Phi}_{(Q,\psi_{2})}$.
@@
By Lemma 1 in the preprint version of our conference paper [BG24b], it immediately follows that $\forall m\in\mathcal{M}^{Q}_{H}:m(v)=m^{\prime}(v)\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q})$. Assuming that the lemma holds for the localized RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, by the definition of $RPS_{1}$ it must also hold that $\forall(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q}):m\models\psi_{1}\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{(Q,\psi_{1})})$ and $\forall(m_{1},\phi_{1})\in\mathcal{C}(p^{\Phi}_{(Q,\psi_{1})}):m_{1}\models\psi_{1}$. Assuming the lemma also holds for $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, from the semantics of the marking-sensitive semi-join and the definition of $RPS_{2}$ it then also follows that $\forall(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q}):m\models\psi_{1}\wedge m\models\psi_{2}\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{(Q,\psi_{2})})$ and $\forall(m_{2},\phi_{2})\in\mathcal{C}(p^{\Phi}_{(Q,\psi_{2})}):m_{2}\models\psi_{2}$. From the semantics of the marking-sensitive semi-join then follows the satisfaction of the Lemma.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $(X^{\Phi},p_{X}^{\Phi})$ a modular extension of the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary configuration for $(X^{\Phi},p_{X}^{\Phi})$. Executing $(X^{\Phi},p_{X}^{\Phi})$ via $O=order^{\Psi}(N^{\Phi})$ then yields a configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ that is consistent for all nodes in $N^{\Phi}$.

### v2

@@ Appendix A Technical Details
For each nested query with the form $(Q^{\prime},\exists(a:Q^{\prime}\rightarrow Q^{\prime\prime},\psi^{\prime}))$, $N^{sat}$ contains the subnets $(N^{\Phi}_{Q^{\prime}},p^{\Phi}_{Q^{\prime}})=localize(Q^{\prime})$ and $(N^{\Phi}_{Q^{\prime\prime}},p^{\Phi}_{Q^{\prime\prime}})=localize(Q^{\prime\prime})$, a marking-sensitive union node $[\cup]^{\Phi}$, a marking-sensitive semi-join $[\ltimes]^{\Phi}$, as well as the request projection structure $RPS^{\infty}_{l}=RPS^{\infty}([\cup]^{\Phi},N^{\Phi}_{Q^{\prime}})$. $N$ contains a RETE subnet $(N_{Q^{\prime}},p_{Q^{\prime}})$ computing matches for $Q^{\prime}$ and a semi-join []delimited-[]left-normal-factor-semidirect-product[$\ltimes$][ ⋉ ].
By Theorem 5 and because the additional dependency on $RPS^{\infty}_{l}$ does not affect the worst case of a fully populated net $(N^{\Phi}_{Q},p^{\Phi}_{Q})$, it holds that $\sum_{n^{\Phi}_{Q^{\prime}}\in V^{N^{\Phi}_{Q^{\prime}}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q^{\prime}})}|m|\leq 7\cdot|\mathcal{C}|_{N_{Q^{\prime}}}|_{e}$. Also, by the semantics of the marking-sensitive semi-join and because the result set for $p^{\Phi}_{Q^{\prime}}$ contains each match for $Q^{\prime}$ at most once by Lemma 12 in [BG24b], it must hold that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\ltimes]^{\Phi})}|m|\leq\sum_{m\in\mathcal{C}(p_{Q^{\prime}})}|m|\leq|\mathcal{C}|_{N_{Q^{\prime}}}|_{e}$.
Any nested query with the form $(Q^{\prime},\exists(a:Q^{\prime}\rightarrow Q^{\prime\prime},\psi^{\prime}))$ must have some nested query $(Q^{\prime\prime},\exists(a:Q^{\prime\prime}\rightarrow Q^{\prime\prime\prime},\psi^{\prime\prime}))$ or $(Q^{\prime\prime},\texttt{true})$ in order for $\psi$ by the definition of nested graph conditions. In either case, $N$ has to contain a corresponding subnet $(N_{Q^{\prime\prime}},p_{Q^{\prime\prime}})$ computing matches for $Q^{\prime\prime}$. It then holds for this subnet $N_{Q^{\prime\prime}}$ by Theorem 5 that $\sum_{n^{\Phi}_{Q^{\prime\prime}}\in V^{N^{\Phi}_{Q^{\prime\prime}}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{Q^{\prime\prime}})}|m|\leq 7\cdot|\mathcal{C}|_{N_{Q^{\prime\prime}}}|_{e}$. It also must hold that $\sum_{(m,\phi)\in\mathcal{C}^{\Phi}([\cup]^{\Phi})}|m|\leq\sum_{m\in\mathcal{C}(p_{Q^{\prime\prime}})}|m|\leq|\mathcal{C}|_{N_{Q^{\prime\prime}}}|_{e}$ and $\sum_{n^{\Phi}\in V^{RPS^{\infty}_{l}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 2\cdot|\mathcal{C}|_{N_{Q^{\prime\prime}}}|_{e}$.
@@
(RETE net localization via $localize^{sat}$ introduces only a constant factor overhead on memory consumption.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ for host graph $H$ and relevant subgraph $H_{p}$ corresponding to $(N,p)$. Assuming that storing a match $m$ requires an amount of memory in $O(|m|)$ and storing an element from $\overline{\mathbb{N}}$ requires an amount of memory in $O(1)$, storing $\mathcal{C}^{\Phi}$ requires memory in $O(|\mathcal{C}|_{e})$.
@@
(Execution time overhead introduced by RETE localization via $localize^{sat}$ depends on query graph size compared to average match size.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a RETE net created via the procedure described in [BG23] for the extended graph query $(Q,\psi)$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}_{0}$ the empty configuration for the marking-sensitive RETE net $(N^{sat},p^{sat})=localize^{sat}(Q,\psi)$ corresponding to $(N,p)$. Executing $(N^{\Phi},p^{\Phi})$ via $execute(order^{\Psi}(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ then takes $O(T\cdot|Q_{max}|)$ steps, with $T=\sum_{n\in V^{N}}|\mathcal{C}(n)|$ and $|Q_{max}|$ the size of the largest graph among $Q$ and graphs in $\psi$.
@@ A.5. Supplementary Lemmata
(RETE net localization via $localize$ introduces an overhead of at most factor 5 on effective configuration size for the production node and at most factor 7 for all other nodes.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net with $Q$ the associated query graph of $p$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}\sum_{(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})}|m|\leq 7\cdot\sum_{n\in V^{N}\setminus\{p\}}\sum_{m\in\mathcal{C}(n)}|m|+5\cdot\sum_{m\in\mathcal{C}(p)}|m|$.
@@
(RETE net localization via $localize$ introduces an overhead of factor 7 on the number of computed matches for the production node and factor 10 for all other nodes.) Let $H$ be an edge-dominated graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net with $Q$ the associated query graph of $p$, $\mathcal{C}$ a consistent configuration for $(N,p)$ for host graph $H$, and $\mathcal{C}^{\Phi}$ a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ for host graph $H$ and relevant subgraph $H_{p}$. It then holds that $\sum_{n^{\Phi}\in V^{N^{\Phi}}}|\mathcal{C}^{\Phi}(n^{\Phi})|\leq 10\sum_{n\in V^{N}\setminus\{p\}}|\mathcal{C}(n)|+7\times|\mathcal{C}(p)|$.
@@
(Consistent configurations for modular extensions of RETE nets localized via $localize^{\Psi}$ yield only matches that satisfy the NGC and all admissible matches with a partial match marked $\infty$ in an extension point.) Let $H$ be a graph, $(Q,\psi)$ a graph query, $(X^{\Phi},p_{X}^{\Phi})$ a modular extension of the localized RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\phi)$, and $\mathcal{C}$ a configuration for $(X^{\Phi},p_{X}^{\Phi})$ that is consistent for all nodes in $V^{N^{\Phi}}$. Furthermore, let $x\in\chi(N^{\Phi})$, $Q_{v}=(\{v\},\emptyset,\emptyset,\emptyset)$ the query subgraph associated with $x$, and $(m^{\prime},\infty)\in\mathcal{C}(x)$ for some match $m^{\prime}:Q_{v}\rightarrow H$. It then holds that $\forall(m,\phi)\in\mathcal{C}(p^{\Phi}):m\models\psi\wedge\forall m\in\mathcal{M}^{Q}_{H}:m(v)=m^{\prime}(v)\wedge m\models\psi\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi})$.
@@
For a nested condition of the form $\psi=\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, $N^{\Phi}$ consists of the localized RETE net for $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, a request projection structure $RPS^{\infty}_{l}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q^{\prime},\psi^{\prime})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q^{\prime},\psi^{\prime})}$.
In the case where $a$ is a partial graph morphism from some subgraph $Q_{p}\subseteq Q$ into $Q^{\prime}$, the argumentation works analogously.
@@
By Lemma 1 in the preprint version of our conference paper [BG24b], it immediately follows that $\forall m\in\mathcal{M}^{Q}_{H}:m(v)=m^{\prime}(v)\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q})$. Assuming that the Lemma holds for the localized RETE net $(N^{\Phi}_{(Q^{\prime},\psi^{\prime})},p^{\Phi}_{(Q^{\prime},\psi^{\prime})})=localize^{\Psi}(Q^{\prime},\psi^{\prime})$, by the definition of $RPS^{\infty}_{l}$ it must also hold that $\forall(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q}),m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m=m^{\prime}\circ a\wedge m^{\prime}\models\psi^{\prime}\Rightarrow(m^{\prime},\infty)\in\mathcal{C}(p^{\Phi}_{(Q^{\prime},\psi^{\prime})})$ and $\forall(m^{\prime},\phi^{\prime})\in\mathcal{C}(p^{\Phi}_{(Q^{\prime},\psi^{\prime})}):m^{\prime}\models\psi^{\prime}$. From the semantics of the marking-sensitive semi-join then follows the satisfaction of the Lemma.
@@
For a nested condition of the form $\psi=\neg\psi^{\prime}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, a request projection structure $RPS^{\infty}_{l}=RPS(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi^{\prime})})$, and a marking-sensitive anti-join node $p^{\Phi}=[\rhd]^{\Phi}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q,\psi^{\prime})}$.
@@
By Lemma 1 in the preprint version of our conference paper [BG24b], it immediately follows that $\forall m\in\mathcal{M}^{Q}_{H}:m(v)=m^{\prime}(v)\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q})$. Assuming that the Lemma holds for the localized RETE net $(N^{\Phi}_{(Q,\psi^{\prime})},p^{\Phi}_{(Q,\psi^{\prime})})=localize^{\Psi}(Q,\psi^{\prime})$, by the definition of $RPS^{\infty}_{l}$ it must also hold that $\forall(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q}):m\models\psi^{\prime}\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{(Q,\psi^{\prime})})$ and $\forall(m,\phi)\in\mathcal{C}(p^{\Phi}_{(Q,\psi^{\prime})}):m\models\psi^{\prime}$. From the semantics of the marking-sensitive anti-join then follows the satisfaction of the Lemma.
@@
For a nested condition of the form $\psi=\psi_{1}\wedge\psi_{2}$, $N^{\Phi}$ consists of the localized RETE net for the plain pattern $Q$, $(N^{\Phi}_{Q},p^{\Phi}_{Q})=localize(Q)$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, a request projection structure $RPS^{\infty}_{1}=RPS^{\infty}(p^{\Phi}_{Q},N^{\Phi}_{(Q,\psi_{1})})$, a marking-sensitive semi-join node $[\ltimes]^{\Phi}_{1}$ with left dependency $p^{\Phi}_{Q}$ and right dependency $p^{\Phi}_{(Q^{\prime},\psi^{\prime})}$, the localized RETE net $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, a request projection structure $RPS^{\infty}_{2}=RPS^{\infty}([\ltimes]^{\Phi}_{1},N^{\Phi}_{(Q,\psi_{2})})$, and a marking-sensitive semi-join node $p^{\Phi}=[\ltimes]^{\Phi}_{2}$ with left dependency $[\ltimes]^{\Phi}_{1}$ and right dependency $p^{\Phi}_{(Q,\psi_{2})}$.
@@
By Lemma 1 in the preprint version of our conference paper [BG24b], it immediately follows that $\forall m\in\mathcal{M}^{Q}_{H}:m(v)=m^{\prime}(v)\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q})$. Assuming that the lemma holds for the localized RETE net $(N^{\Phi}_{(Q,\psi_{1})},p^{\Phi}_{(Q,\psi_{1})})=localize^{\Psi}(Q,\psi_{1})$, by the definition of $RPS^{\infty}_{1}$ it must also hold that $\forall(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q}):m\models\psi_{1}\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{(Q,\psi_{1})})$ and $\forall(m_{1},\phi_{1})\in\mathcal{C}(p^{\Phi}_{(Q,\psi_{1})}):m_{1}\models\psi_{1}$. Assuming the lemma also holds for $(N^{\Phi}_{(Q,\psi_{2})},p^{\Phi}_{(Q,\psi_{2})})=localize^{\Psi}(Q,\psi_{2})$, from the semantics of the marking-sensitive semi-join and the definition of $RPS^{\infty}_{2}$ it then also follows that $\forall(m,\infty)\in\mathcal{C}(p^{\Phi}_{Q}):m\models\psi_{1}\wedge m\models\psi_{2}\Rightarrow(m,\infty)\in\mathcal{C}(p^{\Phi}_{(Q,\psi_{2})})$ and $\forall(m_{2},\phi_{2})\in\mathcal{C}(p^{\Phi}_{(Q,\psi_{2})}):m_{2}\models\psi_{2}$. From the semantics of the marking-sensitive semi-join then follows the satisfaction of the Lemma.
@@
(Execution of modular extensions of localized RETE nets via $order^{\Psi}$ yields consistent configurations for the base RETE net.) Let $H$ be a graph, $H_{p}\subseteq H$, $(Q,\psi)$ a graph query, and $(X^{\Phi},p_{X}^{\Phi})$ a modular extension of the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize^{\Psi}(Q,\psi)$. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary configuration for $(X^{\Phi},p_{X}^{\Phi})$. Executing $(X^{\Phi},p_{X}^{\Phi})$ via $O=order^{\Psi}(N^{\Phi})$ then yields a configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$ that is consistent for all nodes in $N^{\Phi}$.